A 30 kg falling object hits the ground at a speed of 11.5 m/s. If the kinetic energy of the object when it hits the ground is equal to the gravitational potential energy it had before it was dropped, at what height was it dropped from?


1
Expert's answer
2021-06-13T07:09:22-0400
"KE=PE,""\dfrac{1}{2}mv^2=mgh,""h=\dfrac{v^2}{2g}=\dfrac{(11.5\ \dfrac{m}{s})^2}{2\cdot9.8\ \dfrac{m}{s^2}}=6.75\ m."